OpenStudy (anonymous):

Variable y varies directly with x^2, and y = 96 when x = 4 What is the value of y when x = 2? Y =

OpenStudy (anonymous):

I already did some of it but the end is confusing I already did this Y = kx^2 K = y/x^2 K = 96/4^2 But I don't know how to substitute 96 for y and 4 for x

jimthompson5910 (jim_thompson5910):

your next step is to evaluate `96/4^2`

jimthompson5910 (jim_thompson5910):

keep in mind that PEMDAS says you have to do the exponents first, then division comes after (since the E for exponent comes before D for division)

OpenStudy (anonymous):

Okay thanks

jimthompson5910 (jim_thompson5910):

let me know what you get

OpenStudy (anonymous):

I'm still confused I can't solve it @jim_thompson5910

jimthompson5910 (jim_thompson5910):

so you are not able to evaluate what `96/4^2` is equal to?

jimthompson5910 (jim_thompson5910):

does it help if I replaced the `4^2` with `16` ? so `96/4^2` turns into `96/16` which becomes what?

OpenStudy (anonymous):

No i don't know how

OpenStudy (anonymous):

Yeah I'll try again

jimthompson5910 (jim_thompson5910):

what is 96/16 equal to?

OpenStudy (anonymous):

6?

jimthompson5910 (jim_thompson5910):

yes, k = 6

jimthompson5910 (jim_thompson5910):

so `y = k*x^2` turns into `y = 6*x^2`

jimthompson5910 (jim_thompson5910):

now plug `x = 2` into `y = 6*x^2`

OpenStudy (anonymous):

Y = 6*x^2 so y = 24?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

don't forget to show the step `y = 6*2^2`

OpenStudy (anonymous):

Thank you!!!!!!

jimthompson5910 (jim_thompson5910):

you're welcome